# L2R: Learning to Reduce Search Space for Generalizable Neural Routing Solver

### v1

@@ 5.2 Performance Evaluation
Cross-Distribution VRPs
We evaluate L2R’s cross-distribution performance on TSP5K/CVRP5K instances from three distinct distributions: cluster, explosion, and implosion. As shown in Table 5, L2R consistently achieves the best performance among all comparable methods across these distributions. These results further highlight L2R’s robust generalization capabilities.
Benchmark Dataset
We further assess L2R’s generalization performance on instances from CVRPLIB Set-XXL (Arnold et al., 2019) and TSPLIB (Reinelt, 1991). As demonstrated in Table 6, L2R maintains its position as the best-performing model across instances of varying scales, underscoring its practical applicability in real-world scenarios. Detailed results are provided in Appendix F.
5.3 Ablation Study
Effects of Learning-based Reduction
To validate the effectiveness of learning-based reduction in L2R, we train a new model using dynamic distance-based SSR while keeping all other settings unchanged. The results demonstrate that L2R consistently outperforms this alternative with dynamic SSR, particularly in solving cross-distribution instances. Detailed results are provided in Section G.1.
@@ 5.3 Ablation Study
L2R vs. Distance-based SSR with Larger Search Space
We train three dynamic distance-based SSR models with varying search space sizes ($k=50,75,100$) and compare them with our proposed L2R model ($k=20$) on TSP instances of different problem sizes. The results show that L2R, despite its smaller search space, consistently outperforms the distance-based SSR models with larger search spaces across instances of varying scales. Detailed results are available in Section G.2.
6 Conclusion
In this work, we propose a novel RL-based Learning-to-Reduce (L2R) framework for solving large-scale vehicle routing problems. Our approach adaptively selects a small set of candidate nodes at each construction step, enabling efficient search space reduction while maintaining solution quality. Extensive experiments demonstrate that L2R, trained exclusively on 100-node instances from the uniform distribution, achieves remarkable performance on TSP and CVRP instances with up to 1 million nodes from uniform distribution and over 80K nodes from other distributions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Learning-based Reduction Model for CVRPs
@@
Given an instance $S=\{\mathbf{s}_{i}\}_{i=1}^{N+1}$ with node features $\mathbf{s}_{i}\in\mathbb{R}^{d_{x}}$ (i.e., node coordinates and demand), we separate the depot and nodes and feed them into distinct linear layers. The detailed calculation is expressed as
$\mathbf{h}_{i}^{D}=\begin{cases}W_{\text{depot}}^{(e)}[x_{i},\ y_{i}]+\mathbf{b}_{\text{depot}}^{(e)},&i=0\\ W_{\text{node}}^{(e)}[x_{i},\ y_{i},\ \delta_{i}]+\mathbf{b}_{\text{node}}^{(e)},&i=1,\dots,N\end{cases},$ (13)
where $x_{i}$ and $y_{i}$ denote the node coordinates, $\delta_{i}$ is the demand of node $i$, $W_{\text{depot}}^{(e)}\in\mathbb{R}^{2\times d}$ and $\mathbf{b}_{\text{depot}}^{(e)}\in\mathbb{R}^{d}$ are learnable parameters for the depot, and $W_{\text{node}}^{(e)}\in\mathbb{R}^{3\times d}$ and $\mathbf{b}_{\text{node}}^{(e)}\in\mathbb{R}^{d}$ are learnable parameters for the nodes. This process yields a set of embeddings $H^{D}=\{\mathbf{h}_{i}^{D}\}_{i=1}^{N+1}$ for all nodes in the instance $S$.
At the $t$-th step, with a partial solution following Kool et al. (2019) and Kwon et al. (2020), we define a context embedding for the current partial solution:
@@
Appendix B Local Solution Construction Model
@@
B.1 Coordinate Normalization
We introduce a coordinate normalization operation to ensure that each extracted sub-graph $G_{\text{sub}}^{{}^{\prime}}$ adheres to a similar distribution (Fu et al., 2021; Ye et al., 2024; Fang et al., 2024). This operation not only simplifies the input feature space but also significantly enhances the stability and homogeneity of model inputs across different sub-graphs. Following Fu et al. (2021) and Fang et al. (2024), the coordinate transformation is formulated as
@@ B.2 Embedding Layer
Given the converted sub-graph $G_{\text{sub}}^{{}^{\prime\prime}}$, the embedding layer first transforms the coordinates into initial embeddings using a shared linear layer with learnable parameters $[W^{(0)}\in\mathbb{R}^{d_{x}\times d};b^{(0)}\in\mathbb{R}^{d}]$. The embeddings of the $k$ candidate nodes $\mathcal{N}_{t}$ are denoted by $H_{\mathcal{N}}^{(0)}=\{\mathbf{h}_{i}^{(0)}\mid i\in\mathcal{N}_{t}\}\in\mathbb{R}^{k\times d}$. Here, the first node $\pi_{1}$ and the last node $\pi_{t-1}$ are used to represent the current partial solution. Therefore, their initial embeddings require special treatment (Drakulic et al., 2023; Luo et al., 2023). Specifically, additional learnable matrices $W_{1}\in\mathbb{R}^{d\times d}$ and $W_{2}\in\mathbb{R}^{d\times d}$ are applied to $\mathbf{h}_{\pi_{1}}^{(0)}$ and $\mathbf{h}_{\pi_{t-1}}^{(0)}$, respectively. Accordingly, we define the initial graph node embeddings $\widetilde{H}^{(0)}\in\mathbb{R}^{(2+|\mathcal{N}_{t}|)\times d}$ as
@@
For CVRP, due to the demands and capacity constraints, we introduce a specialized treatment for the initial embedding of each node. Specifically, the node demands are normalized as $\delta_{i}=\{\delta/Q_{\text{remain}}\mid i\in\mathcal{N}_{t}\}$, where $Q_{\text{remain}}$ represents the remaining capacity. We separate the coordinates and demands and feed them into distinct linear layers. The detailed calculation is expressed as
@@
B.3 Attention Layer
Inspired by BQ (Drakulic et al., 2023) and LEHD (Luo et al., 2023), our attention layer adopts a heavy decoder structure and consists of two sub-layers: an attention sub-layer and a feed-forward (FF) sub-layer. Both sub-layers incorporate layer normalization (Ba et al., 2016) and skip connections (He et al., 2016).
Let $\widetilde{H}^{(\ell-1)}=[\mathbf{h}_{\pi_{1}}^{(\ell-1)},\ \mathbf{h}_{\pi_{t-1}}^{(\ell-1)},\ H_{\mathcal{N}_{t}}^{(\ell-1)}]$ denote the input to the $\ell$-th attention layer for $\ell=1,\ldots,L$. The outputs for the $i$-th node are computed as follows:
@@ B.3 Attention Layer
where $\mathrm{LN}(\cdot)$ denotes layer normalization (Ba et al., 2016), which mitigates potential value overflows caused by exponential operations; $\mathrm{Attention}$ in Equation 20 represents the adopted attention mechanism (see Appendix C for details); and $\mathrm{FF}(\cdot)$ in Equation 21 corresponds to a fully connected neural network with ReLU activation.
After $L$ attention layers, the final node embeddings $\widetilde{H}^{(L)}=[\mathbf{h}_{\pi_{1}}^{(L)},\ \mathbf{h}_{\pi_{t-1}}^{(L)},\ H_{\mathcal{N}_{t}}^{(L)}]\in\mathbb{R}^{(2+|\mathcal{N}_{t}|)\times d}$ encapsulate the advanced feature representations of the first node, last node, and $k$ candidate nodes.
B.4 Padding
@@
Appendix C Adaptation Attention Free Module
Following Zhou et al. (2024), we implement $\mathrm{Attention}(\cdot)$ using a scale-distance Adaptation Attention Free Module (AAFM) to enhance geometric pattern recognition for routing problems. Given the input $X$, AAFM first transforms it into $Q$, $K$, and $V$ through corresponding linear projection operations:

### v2

@@ A.1 Traveling Salesman Problem (TSP)
A.2 Capacitated Vehicle Routing Problem (CVRP)
A general formulation for CVRP with maximum vehicle number constraints $\epsilon_{m}$[41, 34] is introduced in this section. Specifically, given a complete graph $G=(\Omega,E)$, binary variables $z_{ij}\in\{0,1\}$ are defined to represent whether a vehicle traverses arc $(i,j)$ in the solution. $c_{i}$ and $\delta_{i}$ denote the remaining vehicle capacity after visiting customer $i$ and the demand of customer $i$, respectively. The maximum capacity of each vehicle is denoted as $C$, respectively. Given these variables, the formulation is given as follows.
minimize $\sum_{i\in\Omega,j\in\Omega,i\neq j}{d_{ij}z_{ij}}$ (16)
subject to $\sum_{j\in\Omega,i\neq j}x_{ij}=1,\quad\forall i\in\Omega\backslash\{0\}$ (17)
$\sum_{i\in\Omega,i\neq j}x_{ij}=1,\quad\forall j\in\Omega\backslash\{0\}$ (18)
$c_{j}=\begin{cases}c_{i}-\delta_{j},&\text{if }z_{ij}=1,\quad\forall(i,j)\in\Omega\\ c_{j},&\text{if }z_{ij}=0,\quad\forall(i,j)\in\Omega\end{cases}$ (19)
$c_{j}-c_{i}\geq\delta_{j}-C(1-z_{ij}),\quad\forall i,j\in\Omega\backslash\{0\},i\neq j,d_{i}+d_{j}\leq C$ (20)
$0\leq c_{i}\leq C-\delta_{i},\quad i\in\Omega\backslash\{0\}$ (21)
$\sum_{j\in\Omega\backslash\{0\}}z_{j0}=\epsilon$ (22)
$\sum_{i\in\Omega\backslash\{0\}}z_{i0}=\epsilon$ (23)
$\epsilon\leq\epsilon_{m}$ (24)
$z_{ij}\in\{0,1\},\quad\forall i,j\in\Omega,i\neq j$ (25)
where Equation 16 is objective function, $d_{ij}$ is the Euclidean distance from customer $i$ to customer $j$. Equation 17 and Equation 18 ensure that each customer is visited exactly once. Equation 19 is the updating process of remaining vehicle capacity and Equation 20--Equation 21 define the capacity and connectivity constraints. Equation 22--Equation 24 ensure that the number of vehicles departing from the depot equals the total number returning to the depot, i.e., each vehicle must start and terminate its route at the depot.
Appendix B Related Work
B.1 NCO without Search Space Reduction
Most neural combinatorial optimization (NCO) models are trained on small-scale instances (e.g., 100 nodes) without search space reduction and achieve strong performance on instances of similar size. However, their effectiveness significantly diminishes when applied to larger instances (e.g., those exceeding 1,000 nodes) [42, 43, 44, 45, 46]. Some approaches incorporate additional search procedures, such as 2-opt [47] and active search [14, 48], to address this limitation. While these techniques can improve solution quality, they are still computationally expensive for large-scale instances. Another line of research focuses on training NCO models directly on larger-scale instances (e.g., up to 500 nodes) to enhance generalization [27, 38, 17]. However, this approach incurs prohibitive computational costs due to the exponentially growing search space. Alternatively, some methods simplify large-scale VRPs through decomposition policies [49, 50, 34, 51, 52]. Although effective, these methods often overlook the dependency between decomposition policies and subsequent solvers, resulting in suboptimal solutions. Moreover, the reliance on expert-designed policies limits their practicality for real-world applications. In addition, recent learning-based advancements also show promising results for non-routing problems, such as Job Shop Scheduling Problems [53, 54] and other graph problems (e.g., Maximum Independent Set and Maximum Clique) [55]. These approaches have proven particularly effective in dynamic environments where traditional methods face challenges.
B.2 NCO with Static Search Space Reduction
To address the scalability challenges of NCO methods, static search space reduction (SSR) has been proposed as a computationally efficient approach. These methods perform one-time pruning at the beginning of the optimization process, significantly reducing the problem size. For example, Sun et al. [21] develop a static problem reduction technique to eliminate unpromising edges in large-scale TSP instances. Recently, heatmap-based approaches have gained popularity for solving large-scale TSPs, where models are trained to predict the probability of each edge belonging to the optimal solution. To handle large-scale instances, these methods often incorporate graph sparsification [23, 24, 25, 56] or pruning strategies [57, 22, 58] to reduce the search space. While static SSR is computationally efficient, it typically requires additional well-designed search procedures (e.g., Monte Carlo Tree Search for TSP) to achieve high-quality solutions, which might be more important for the optimization process [59].
B.3 NCO with Dynamic Search Space Reduction
Dynamic search space reduction (SSR) has also been explored as a promising approach to address the scalability challenges of NCO methods. These methods adaptively prune the search space to a small set of candidate nodes at each construction step, typically based on the distance to the last visited node. The final node selection can be guided by either the original policy augmented with auxiliary distance information [27] or a well-designed local policy [26]. Additionally, recent work by Fang et al. [28] and Drakulic et al. [10] directly selects the next node from the candidate set using NCO models. While dynamic SSR can more efficiently reduce the search space for constructive NCO, its reliance on distance-based node selection is inappropriate for large-scale instances, especially those from non-uniform distributions.
@@ B.3 NCO with Dynamic Search Space Reduction
Appendix C Further Analysis of Distance-based Search Space Reduction
C.1 Impact on Optimality Gap
In this subsection, we provide detailed solution visualizations across TSPLIB instances of varying scales under different levels of search space reduction ($k$). As shown in Figure 5, this degradation is primarily attributed to the elimination of non-local node selections that are crucial for optimal routes, and the cumulative effect of over-pruning during the solution construction process leads to significant deviations from the optimal route in subsequent node selections, ultimately compromising the overall solution quality obtained by D-SSR.
C.2 Impacts on Constructive NCOs
In this subsection, regarding the impacts of distance-based search space reduction across diverse constructive NCO architectures, we conduct further experimental analyses, which mainly include: (1) the impact for constructive NCO methods with a heavy decoder, (2) the impact for constructive NCO methods with a heavy encoder, and (3) the impact for the retrained constructive NCO method.
(a) LEHD with Light Encoder and Heavy Decoder.
(b) POMO with Heavy Encoder and Light Decoder.
Figure 4: The effect of distance-based search space reduction on the gap in diverse constructive NCO architectures (Left: LEHD [9]; Right: POMO [18]). Here $\Delta Gap=Gap_{\text{reduced}}-Gap_{\text{original}}$.
(a) d1291 without D-SSR   (Scale: 1291, Gap: 0.00%)
(b) d1291 with D-SSR of $k=10$   (Scale: 1291, Gap: 26.71%)
(c) d1291 with D-SSR of $k=20$   (Scale: 1291, Gap: 9.59%)
(d) d1655 without D-SSR   (Scale: 1655, Gap: 0.00%)
(e) d1655 with D-SSR of $k=10$   (Scale: 1655, Gap: 20.41%)
(f) d1655 with D-SSR of $k=20$   (Scale: 1655, Gap: 8.74%)
(g) rl1889 without D-SSR   (Scale: 1889, Gap: 0.00%)
(h) rl1889 with D-SSR of $k=10$   (Scale: 1889, Gap: 27.94%)
(i) rl1889 with D-SSR of $k=20$   (Scale: 1889, Gap: 8.45%)
(j) usa13509 without D-SSR   (Scale: 13509, Gap: 0.01%)
(k) usa13509 with D-SSR of $k=10$   (Scale:13509, Gap:20.06%)
(l) usa13509 with D-SSR of $k=20$   (Scale: 13509, Gap: 5.87%)
Figure 5: The solution visualizations of TSPLIB instances under different levels of search space. All solutions are obtained by the LKH-3 algorithm [3].
C.2.1 Impact on Pretrained Constructive NCO
Impact on Constructive NCO with Heavy Decoder
To evaluate the impact of search space reduction on NCO performance, we use the well-known LEHD method [9] as a representative example, restricting the search space to the $k$ nearest nodes from the last visited node at each construction step. Without loss of generality, we use the LEHD-greedy without any RRC strategy. We quantify the impact by measuring the performance gap $\Delta Gap=Gap_{\text{reduced}}-Gap_{\text{original}}$, defined as the difference between the optimality gap obtained by LEHD with a reduced search space ($Gap_{\text{reduced}}$) and the original LEHD without search space reduction ($Gap_{\text{original}}$).
The results in Figure 4(a) illustrate the performance on TSP instances of varying sizes under different search space reduction levels ($k$). A key observation is the existence of a critical threshold $k^{*}$ for each problem size, beyond which LEHD with $k\geq k^{*}$ achieves a $0\%$ performance gap with the original LEHD, indicating no loss in solution quality with SSR. However, these critical thresholds vary significantly across problem sizes, and using a small $k<k^{*}$ leads to substantial performance degradation.
Impact on Constructive NCO with Heavy Encoder
In addition to SL-based LEHD with a heavy decoder architecture, we also evaluate the RL-based POMO model featuring a heavy encoder architecture, and the instance augmentation technology is not used in our experiment. As shown in Figure 4(b), our experimental results demonstrate a consistent phenomenon with the SL-based LEHD case: The critical thresholds $k^{*}$ exhibit significant variations across different problem sizes, and using a small $k<k^{*}$ results in considerable performance deterioration.
C.2.2 Impact on Retrained Constructive NCO
Considering that the inconsistency between training and inference strategies may introduce a deviation between the model’s learned perception and its actual decision-making process, we retrain three reduced variants of LEHD with different search space reduction configurations. As illustrated in Table 8, while reducing the search space significantly decreases inference time, employing an excessively small $k$ based on pairwise distance information results in substantial performance deterioration.
Table 8: The effect of distance-based search space reduction on the gap in the LEHD model. Note that all LEHD models with search space reduction are retrained with the same specifications.
Method TSP100 TSP200 TSP500 TSP1000
Gap Time Gap Time Gap Time Gap Time
LKH3 0.00% 56m 0.00% 4m 0.00% 32m 0.00% 8.2h
LEHD w/ $k=20$ 3.10% 10.7s 5.21% 1.0s 7.95% 2.1s 11.09% 3.7s
LEHD w/ $k=30$ 1.71% 15s 3.03% 1.1s 6.01% 2.2s 9.44% 4s
LEHD w/ $k=50$ 0.87% 18.7s 1.74% 1.28s 4.03% 2.6s 7.41% 4.9s
LEHD w/o SSR 0.58% 24s 0.86% 3s 1.56% 18s 3.17% 1.6m
These findings highlight the potential of search space reduction while underscoring the need for advanced reduction methods capable of efficiently handling large-scale problems with a small $k$.
Appendix D Compatibility with Adaptation Bias
As defined in Equation 6, the adaptation bias between each feasible node $i\in\mathcal{A}_{t}$ and the current node $\pi_{t-1}$ (selected by the local construction model) is given by $a_{t-1,i}^{R}=-\alpha\cdot\log_{2}{N}\cdot d_{t-1,i}$, following Zhou et al. [17]. Here, $N$ denotes the total number of nodes (i.e., problem size), and $\alpha>0$ is a learnable parameter with a default value of 1 that enables the model to adaptively learn information-specific weights. This function incorporates both distance $d_{t-1,i}$ and scale information $\log_{2}{N}$ to better capture varying geometric variations with different scales. Accurately identifying candidate nodes within a large search space remains challenging when relying solely on a lightweight network architecture. Unlike existing distance-based reduction methods, our approach addresses the trade-off between learning difficulty and solution quality by incorporating a distance-assisted reduction model to enhance neighborhood selection performance.
Additionally, we extend this adaptation bias to the compatibility computation in the local construction model. In Equation 10, the bias between each candidate node $i\in\mathcal{N}_{t}$ and the current node $\pi_{t-1}$ is defined as $a_{t-1,i}^{L}=-\alpha\cdot\log_{2}{|\mathcal{N}_{t}|}\cdot d_{t-1,i}$. Here, $|\mathcal{N}_{t}|$ denotes the total number of candidate nodes.
To evaluate the effectiveness of two adaptation biases, we conduct an ablation study with different component configurations, and the detailed results are provided in Section L.2.
Appendix E Learning-based Reduction Model for CVRPs
@@
Given an instance $S=\{\mathbf{s}_{i}\}_{i=1}^{N+1}$ with node features $\mathbf{s}_{i}\in\mathbb{R}^{d_{x}}$ (i.e., node coordinates and demand), we feed depot and nodes into a unified linear layer. The detailed calculation is expressed as
$\mathbf{h}_{i}^{D}=W^{(e)}[x_{i},\ y_{i},\ \delta_{i}]+\mathbf{b}^{(e)},\ i=1,\dots,N+1,$ (26)
where $x_{i}$ and $y_{i}$ denote the node coordinates, $\delta_{i}$ is the demand of node $i$, we regard the demand of depot as 0 (i.e., $\delta_{0}=0$), $W^{(e)}\in\mathbb{R}^{3\times d}$ and $\mathbf{b}^{(e)}\in\mathbb{R}^{d}$ are learnable parameters. This process yields a set of embeddings $H^{D}=\{\mathbf{h}_{i}^{D}\}_{i=1}^{N+1}$ for all nodes in the instance $S$. At the $t$-th step, with a partial solution following Kwon et al. [18], we define a context embedding for the current partial solution:
@@
Appendix F Local Solution Construction Model
@@
F.1 Coordinate Normalization
We introduce a coordinate normalization operation to ensure that each extracted sub-graph $G_{\text{sub}}^{{}^{\prime}}$ adheres to a similar distribution [22, 39, 28]. This operation not only simplifies the input feature space but also significantly enhances the stability and homogeneity of model inputs across different sub-graphs. Following Fu et al. [22] and Fang et al. [28], the coordinate transformation is formulated as
@@ F.2 Embedding Layer
Given the converted sub-graph $G_{\text{sub}}^{{}^{\prime\prime}}$, the embedding layer first transforms the coordinates into initial embeddings using a shared linear layer with learnable parameters $[W^{(0)}\in\mathbb{R}^{d_{x}\times d};b^{(0)}\in\mathbb{R}^{d}]$. The embeddings of the $k$ candidate nodes $\mathcal{N}_{t}$ are denoted by $H_{\mathcal{N}}^{(0)}=\{\mathbf{h}_{i}^{(0)}\mid i\in\mathcal{N}_{t}\}\in\mathbb{R}^{k\times d}$. Here, the first node $\pi_{1}$ and the last node $\pi_{t-1}$ are used to represent the current partial solution. Therefore, their initial embeddings require special treatment [10, 9]. Specifically, additional learnable matrices $W_{1}\in\mathbb{R}^{d\times d}$ and $W_{2}\in\mathbb{R}^{d\times d}$ are applied to $\mathbf{h}_{\pi_{1}}^{(0)}$ and $\mathbf{h}_{\pi_{t-1}}^{(0)}$, respectively. Accordingly, we define the initial graph node embeddings $\widetilde{H}^{(0)}\in\mathbb{R}^{(2+|\mathcal{N}_{t}|)\times d}$ as
@@
For CVRP, due to the demands and capacity constraints, we introduce a specialized treatment for the initial embedding of each node. Specifically, the node demands are normalized as $\delta_{i}=\{\delta_{i}/Q_{\text{remain}}\mid i\in\mathcal{N}_{t}\}$, where $Q_{\text{remain}}$ represents the remaining capacity. We separate the coordinates and demands and feed them into distinct linear layers. The detailed calculation is expressed as
@@
F.3 Attention Layer
Inspired by BQ [10] and LEHD [9], our attention layer consists of two sub-layers: an attention sub-layer and a feed-forward (FF) sub-layer. Both sub-layers incorporate Layer Normalization [30] and skip connections [31].
Let $\widetilde{H}^{(\ell-1)}=[\mathbf{h}_{\pi_{1}}^{(\ell-1)},\ \mathbf{h}_{\pi_{t-1}}^{(\ell-1)},\ H_{\mathcal{N}_{t}}^{(\ell-1)}]$ denote the input to the $\ell$-th attention layer for $\ell=1,\ldots,M$. The outputs for the $i$-th node are computed as follows:
@@ F.3 Attention Layer
where $\mathrm{LN}(\cdot)$ denotes layer normalization [30], which mitigates potential value overflows caused by exponential operations; $\mathrm{Attention}$ in Equation 33 represents the adopted attention mechanism (see Appendix G for details), and $\mathrm{FF}(\cdot)$ in Equation 34 corresponds to a fully connected neural network with ReLU activation.
After $M$ attention layers, the final node embeddings $\widetilde{H}^{(M)}=[\mathbf{h}_{\pi_{1}}^{(M)},\ \mathbf{h}_{\pi_{t-1}}^{(M)},\ H_{\mathcal{N}_{t}}^{(M)}]\in\mathbb{R}^{(2+|\mathcal{N}_{t}|)\times d}$ encapsulate the advanced feature representations of the first node, last node, and $k$ candidate nodes.
F.4 Padding
@@
Appendix G Adaptation Attention Free Module
Following Zhou et al. [17], we implement $\mathrm{Attention}(\cdot)$ using a scale-distance adopted adaptation attention free module (AAFM) to enhance geometric pattern recognition for routing problems. Given the input $X$, AAFM first transforms it into $Q$, $K$, and $V$ through corresponding linear projection operations: